Let a,b, c be integers. Prove that if a|b and a|(b+c) then a|c.
1
Expert's answer
2012-04-03T12:03:24-0400
Let a=14, b=2 and c=5. Obviously, a|b cause 14/2 = 7 and a|(b+c) cause 14/(2+5) = 14/7 = 2. But a isn't divisible by c cause 14 isn't divisible by 5. So, the statement is incorrect.